A dog is at a location 12.1 m behind its owner (x= -12.1m). The owner throws the ball 3m ahead of themselves for their dog to chase (x=3m). The dog is able to run with a velocity of 5 m/s. (A) Where is the dog after 1.6 s of running? (B) How long does it take for the dog to reach the ball?
1
Expert's answer
2020-10-19T13:20:16-0400

The position of the dog is given by the following:


"x= x_0 + vt"

where "x_0 = -12.1 m" is the intial position and "v = 5m\/s". Thus, after "t = 1.6s" the dog will be in:


"x = -12.1m + 5m/s\cdot 1.6s = -4.1m"

The distance between the dog and the ball is:


"d = 3m-(-12.1m) = 15.1m"

Moving with the velocity "v = 5m\/s" the dog covers this distance in:


"t = \dfrac{d}{v} = \dfrac{15.1m}{5m/s} =3.02s"

Answer. a) -4.1m, b) 3.02 s.